# Measuring Bias or Measuring the Task: Understanding the Brittle Nature of LLM Gender Biases

### v1

@@ Abstract
As LLMs are increasingly applied in socially impactful settings, concerns about gender bias have prompted growing efforts both to measure and mitigate such bias. These efforts often rely on evaluation tasks that differ from natural language distributions, as they typically involve carefully constructed task prompts that overtly or covertly signal the presence of gender bias-related content. In this paper, we examine how signaling the evaluative purpose of a task impacts measured gender bias in LLMs. Concretely, we test models under prompt conditions that (1) make the testing context salient, and (2) make gender-focused content salient. We then assess prompt sensitivity across four task formats with both token-probability and discrete-choice metrics. We find that even minor prompt changes can substantially alter bias outcomes, sometimes reversing their direction entirely. Discrete-choice metrics further tend to amplify bias relative to probabilistic measures. These findings do not only highlight the brittleness of LLM gender bias evaluations but open a new puzzle for the NLP benchmarking and development community: To what extent can well-controlled testing designs trigger LLM “testing mode” performance, and what does this mean for the ecological validity of future benchmarks.
@@ 1 Introduction
As Large Language Models (LLMs) are increasingly integrated into critical applications such as recruitment Gan et al. (2024); Times (2024), education Wikipedia contributors (2025); Gan et al. (2023); Dan et al. (2023), and healthcare Wang et al. (2023), concerns over fairness and bias mitigation have gained prominence Valkanova and Yordanov (2024); Warr et al. (2024); Haltaufderheide and Ranisch (2024). Gender bias within these models, if unaddressed, can perpetuate stereotypes and reinforce systemic inequalities Cheng et al. (2023); Kotek et al. (2023). Addressing this issue requires a deep understanding of how, when, and to what extent bias in LLMs emerges.
In order to improve our ability to quantify gender bias, many efforts have focused on developing scenarios that lead models into recommending actions (e.g., rejecting vs. accepting job applications An et al. (2024)) or making linguistic choices (e.g., associating job titles with pronouns Kotek et al. (2023); Dong et al. (2024)) which can then be interpreted in terms of gender bias. However, model benchmarking as a whole constantly plays a game of catch-up: as soon as a new scenario for quantifying gender bias is posed, model development improves upon the benchmark but not necessarily on the more general issue (see, e.g., Kiela et al. (2021) for a broader discussion). [^1] Goodhart’s Law brings it to the point: “When a measure becomes a target, it ceases to be a good measure.” Goodhart (1975) While the testing scenarios are increasingly diverse, they often still either evoke the common evaluation task setup more broadly or introduce a highly gendered context in particular to elicit testable behavior. Benchmarks such as WinoBias Zhao et al. (2018), StereoSet Nadeem et al. (2021), and P-AT Onorati et al. (2023) exemplify this style of instructional or gender-salient framing. [^2] P-AT includes prompts such as “Determine whether an expert on X is more likely to be male or female”; StereoSet uses sentence stems like “Girls tend to be more ___ than boys”; and WinoBias presents multiple-choice options such as “A. he    B. she    C. both.”
@@
There are reasons to believe that each of them might matter, since LLMs are at their core designed to pick up on language distribution shifts and replicate the context-specific linguistic signal. If LLMs have learned how gender bias presents differently in common evaluation task contexts (which are largely available online and presumably in training data), then we might expect their behavior to shift. Similarly, it seems reasonable to assume that linguistic contexts that explicitly discuss gender will be associated with gender representations that are distinct from the distribution in common pretraining data, and should therefore result in distinct bias behavior. The resulting model behavior would then resemble a unique performance during test time (i.e., testing mode behavior) that is not representative of behavior outside of benchmarks. Note that we don’t claim that a testing mode is a distinct internal model state, but a robust behavioral shift induced by prompts that have general features of evaluation benchmarks. [^3] Parallels can be drawn to task demand characteristics in psychology Banaji and Hardin (1996); Greenwald et al. (1998); Oakhill et al. (2005), which we will further discuss in Section 6.
@@ 2 Related Work
Before turning to related work on (1) methods for measuring bias in LLMs and (2) the impact of prompt sensitivity in evaluation, we first clarify our terminology. In this work, we investigate to what extent models display distinct gender bias behavior when they face common (bias) evaluation prompts.
There are two potential prompt formats used in gender bias literature which we aim to disambiguate. Recent work uses instruction-following prompts, where the scenario is framed as a task and metrics quantify bias in instruction-tuned model’s response patterns (e.g., “Based on this CV, which job would you recommend?” Bai et al. (2025)). More traditional setups use “pure” language modeling where next-token prediction is conditioned on the previous context tokens (e.g., “The doctor talked to the patient” Caliskan et al. (2017); Bolukbasi et al. (2016)) or a masked out intermediate token (e.g., “People who say X are” Nozza et al. (2021); Nadeem et al. (2021)). For the purpose of this paper, we use the term prompt as an umbrella term to capture the notion of conditioning a model on prior context to quantify subsequent model behavior. In our experiments, we use the latter setup, since we specifically contrast results with a no-instruction condition.
@@ 2.2 Prompt Variations Highly Affect Model Behavior
Our work builds on much prior work which has documented how seemingly inocuous prompt variations can lead to vastly distinct LLM behaviors  Hu and Levy (2023); Dong et al. (2023); Bai et al. (2025); Sclar et al. (2023). For example,  Shah et al. (2024) introduce POSIT, a Prompt Sensitivity Index quantifying how output distributions change under intent-preserving prompt rewrites. They find that prompt sensitivity persists across model scales and alignment levels, revealing fundamental brittleness. Similarly, Shin et al. (2023) show that simple formatting changes can cause up to 76-point swings in few-shot classification accuracy, highlighting the need for more stable and controlled evaluation pipelines.
@@ 4.2 Gender Inference Metrics
Figure 1: Overall Pronoun Shift Results. Each violin plot shows the distribution of model-level sensitivity scores across all evaluated attributes. The black line indicates the mean sensitivity score, with vertical bars denoting the 95% confidence interval. Wider sections of the violin reflect more frequent sensitivity values.
@@ 4.3 Prompt Sensitivity Evaluation
APD ranges from 0 (identical distributions) to 1 (fully divergent), meaning that the score is zero when the model output distribution doesn’t change between the prompt conditions and one if this change is maximal. This serves as the basis for the two sensitivity scores: the Gender Salience Effect Score and the Instruction Presence Effect Score. The two only vary in the prompt condition we’re marginalizing over.
@@
We compute the sensitivity of the tested LLMs to the prompt conditions in three stages: (1) We compute the Absolute Proportion Difference between matched prompt variants (e.g., $\text{Gender}^{+}\text{Instr}^{+}$ vs. $\text{Gender}^{+}\text{Instr}^{-}$ ) at the attribute level. (2) We average Absolute Proportion Difference Scores across all attributes to obtain Gender Salience Effect and Instruction Presence Effect per task. (3) We average GenEffect and InstrEffect Scores across tasks, representing overall sensitivity to prompt structure.
@@ 5.1 Pronoun Choices Shift when Gender Bias Evaluation is Salient
Figure 1 shows the distribution of overall prompt sensitivity scores across models. If models were insensitive to the prompt conditions, they would assign the same pronoun in a given scenario, resulting in a sensitivity score of 0. If they show maximally distinct behavior in their gender assignment across conditions, the sensitivity score would be 1.
@@
We observe distinct patterns in how individual models respond to each framing dimension. In the Gender Salience condition (Figure 2a), most models exhibit moderate to high sensitivity, with average scores largely around 0.6. This indicates that when the gender-inference nature of the task is made explicit, models frequently adjust their pronoun outputs. A notable exception is Meta-Llama-3.1-8B, which shows low sensitivity when the prompt primes for gender-related concepts, suggesting a relative insensitivity compared to the other models. This effect appears to drive that Meta-Llama-3.1-8B is an outlier in the overall pronoun shift results (Figure 1)
@@
Figure 3: Pronoun-Specific Shift Probabilities across Models. Each bar represents the mean shift in token probability for a given pronoun—he, she, or they—across all prompt conditions and attributes. The shift is computed using relative differences in pronoun probabilities between paired prompt conditions, rather than absolute differences. Error bars indicate 95% confidence intervals, showing variability across attributes.
@@ 5.3 Effects on the Pronoun-Level
While the previous results indicate that using prompts with common bias evaluation setups change model behavior, it leaves open whether these changes in model behavior are interpretable. Prior work has shown that LLMs often default to assigning male gender when the context is ambiguous Kotek et al. (2023); Dong et al. (2024); Kaneko et al. (2024); Tang et al. (2024). Based on the reasoning that LLMs might learn fairer behavior particularly in evaluation settings, we predict that generally underrepresented genders (“she”) and neutral pronouns (singular “they”) should show an increase in assignment in testing scenarios, in contrast to generally overrepresented genders (“he”). The results are shown in  Figure 3. Pronouns that have a sensitivity of zero, don’t change in distribution with varying prompts. Pronouns with a sensitivity < 0 are likely to disappear when the prompt saliently signals gender evaluation and pronouns with a sensitivity > 0 are assigned higher preference.
@@ 5.4 Effects on the Metric-Level
To ensure valid and interpretable data for the proportion of choice analysis, we filter out task-condition pairs in which the model consistently fails to generate pronouns. Specifically, if more than 60% of outputs in a given (model, task, condition) combination contain non-pronoun completions, the setting is considered over-capacity and excluded from analysis. Phi-3-small-128k-instruct and Mistral-Small-Instruct-2409, the smallest models in our evaluation, exhibit the highest number of exclusions, suggesting that potentially limited capacity may impair their ability to provide relevant responses (i.e., outputs containing gendered pronouns or descriptors). Expectedly, instruction-absent conditions were especially noisy in their output but are sufficiently present across models and tasks to allow for an aggregated analysis. We provide all details on the exclusions in Appendix A.
@@ 6 Discussion
Finally, our results have implications for prompt design as intervention. Prompts that foreground gender concepts can shift model outputs in ways that align with fairness goals. This suggests that strategically framed prompts could serve as lightweight mechanisms to influence LLM behavior in practice —though we must be careful not to mistake prompt compliance for true debiasing.
@@
Beyond these observations, our work recommends several directions for more robust evaluation design. Conceptually, this challenge is reminiscent of the problem of task demand characteristics in psychology Banaji and Hardin (1996); Greenwald et al. (1998); Oakhill et al. (2005), where participants adjust their behavior once they infer the purpose of a study. Decades of research has brought forth strategies to minimize task demand in people, and similar approaches may help benchmarks better reveal persistent model tendencies. For example, future benchmarks could explore the benefit of using filler items, refocus on implicit measures (what are reaction time analyses for human studies), or introduce secondary tasks to reduce explicit test awareness. At a more practical level, benchmarks should rely on diverse prompt sets rather than isolated items, report sensitivity ranges to highlight framing effects, and avoid instruction-heavy setups unless fairness alignment is the explicit goal. Together, these practices may help ensure that benchmark outcomes reflect underlying model biases rather than prompt-induced compliance.
@@
We validate the promise of practically incorporating these insights using two recently proposed gender bias benchmarks Dong et al. (2024); Onorati et al. (2023). After replicating their findings, we adapted their prompts to increase the salience of the gender testing variable. (We report all data and implementational details in Appendix B.) In line with our previous results, we find that for both benchmarks and across tested models gender bias scores significantly shift, sometimes even reversing the previously attested bias trend (see Figure 5 for a summary of the main observations). These results emphasize the brittle nature of prompt-based model behavior overall and how gender associations within the task can fundamentally alter gender bias behavior —maybe sometimes even for the better.
@@
Figure 5: Pronoun and Bias Score Shift in Replicate Study One. (a) shows overall pronoun shift; (b) shows change in Gender Attribute Score (GAS). Both reflect aggregate results across all models after prompt modifications. The red line and arrow in (b) indicate the direction and magnitude of the GAS change.

### v3

@@ Abstract
As LLMs are increasingly applied in socially impactful settings, concerns about gender bias have prompted growing efforts both to measure and mitigate such bias. These efforts often rely on evaluation tasks that differ from natural language distributions, as they typically involve carefully constructed task prompts that overtly or covertly signal the presence of gender bias-related content. In this paper, we examine how signaling the evaluative purpose of a task impacts measured gender bias in LLMs. Concretely, we test models under prompt conditions that (1) make the testing context salient, and (2) make gender-focused content salient. We then assess prompt sensitivity across four task formats with both token-probability and discrete-choice metrics. We find that prompts that more clearly align with (gender bias) evaluation framing elicit distinct gender output distributions compared to less evaluation-framed prompts. Discrete-choice metrics further tend to amplify bias relative to probabilistic measures. These findings do not only highlight the brittleness of LLM gender bias evaluations but open a new puzzle for the NLP benchmarking and development community: To what extent can well-controlled testing designs trigger LLM “testing mode” performance, and what does this mean for the ecological validity of future benchmarks.
@@ 1 Introduction
As Large Language Models (LLMs) are increasingly integrated into critical applications such as recruitment Gan et al. (2024), education Wikipedia contributors (2025); Gan et al. (2023); Dan et al. (2024), and healthcare Wang et al. (2023), concerns over fairness and bias mitigation have gained prominence Valkanova and Yordanov (2024); Warr et al. (2024); Haltaufderheide and Ranisch (2024). Gender bias within these models, if unaddressed, can perpetuate stereotypes and reinforce systemic inequalities Cheng et al. (2023); Kotek et al. (2023). Addressing this issue requires a deep understanding of how, when, and to what extent bias in LLMs emerges.
In order to improve our ability to quantify gender bias, many efforts have focused on developing scenarios that lead models into recommending actions (e.g., rejecting vs. accepting job applications An et al. (2024)) or making linguistic choices (e.g., associating job titles with pronouns Kotek et al. (2023); Dong et al. (2024)) which can then be interpreted in terms of gender bias. However, model benchmarking as a whole constantly plays a game of catch-up: as soon as a new scenario for quantifying gender bias is posed, model development improves upon the benchmark but not necessarily on the more general issue (see, e.g., Kiela et al. (2021) for a broader discussion). [^1] Marilyn Strathern’s reframing of Goodhart’s Law brings it to the point: “When a measure becomes a target, it ceases to be a good measure.” Goodhart (1984); Strathern (1997) While the testing scenarios are increasingly diverse, they often still either evoke the common evaluation task setup more broadly or introduce a highly gendered context in particular to elicit testable behavior. Benchmarks such as WinoBias Zhao et al. (2018), StereoSet Nadeem et al. (2021), and P-AT Onorati et al. (2023) exemplify this style of instructional or gender-salient framing. [^2] Examples from P-AT prompts: “Determine whether an expert on X is more likely to be male or female”; StereoSet sentence stems: “Girls tend to be more ___ than boys”; WinoBias multiple-choice options: “A. he    B. she    C. both.”
@@
There are reasons to believe that each of them might matter, since LLMs are at their core designed to pick up on language distribution shifts and replicate the context-specific linguistic signal. If LLMs have learned how gender bias presents differently in common evaluation task contexts (which are largely available online and presumably in training data), then we might expect their behavior to shift. Similarly, it seems reasonable to assume that linguistic contexts that explicitly discuss gender will be associated with gender representations that are distinct from the distribution in common pretraining data, and should therefore result in distinct bias behavior. The resulting model behavior would then resemble a unique performance during test time (i.e., testing mode behavior) that is not generalizable to all behavior outside of benchmarks. Note that we don’t claim that a testing mode is a distinct internal model state, but a robust behavioral shift induced by prompts that have general features of evaluation benchmarks. [^3] Parallels can be drawn to task demand characteristics in psychology Banaji and Hardin (1996); Greenwald et al. (1998); Oakhill et al. (2005), which we further discuss in Section 6.
@@ 2 Related Work
Before turning to related work on (1) methods for measuring bias in LLMs and (2) the impact of prompt sensitivity in evaluation, we first clarify our terminology. There are two prompt formats used in the gender bias literature which we aim to disambiguate. Recent work uses instruction-following prompts, where the scenario is framed as a task and metrics quantify bias in the instruction-tuned model’s response patterns (e.g., “Based on this CV, which job would you recommend?” Bai et al. (2025)). More traditional setups use “pure” language modeling where next-token prediction is conditioned on the previous context tokens (e.g., “The doctor talked to the patient” Caliskan et al. (2017); Bolukbasi et al. (2016)) or a masked out intermediate token (e.g., “People who say X are” Nozza et al. (2021); Nadeem et al. (2021)). For the purpose of this paper, we use prompt as an umbrella term to capture the notion of conditioning a model on any prior context to quantify subsequent model behavior. In our experiments, we use the latter setup, since we specifically contrast results with a no-instruction condition.
@@ 2.2 Prompt Variations Highly Affect Model Behavior
Our work builds on much prior work which has documented how seemingly innocuous prompt variations can lead to vastly distinct LLM behaviors  Hu and Levy (2023); Dong et al. (2023); Bai et al. (2025); Sclar et al. (2024). For example,  Chatterjee et al. (2024) introduce POSIX, a Prompt Sensitivity Index quantifying how output distributions change under intent-preserving prompt rewrites. They find that prompt sensitivity persists across model variants and scales, revealing fundamental brittleness. Similarly, Sclar et al. (2024) show that simple formatting changes can cause up to 76-point swings in few-shot classification accuracy, highlighting the need for more stable evaluation pipelines.
@@ 4.2 Gender Inference Metrics
Figure 1: Probability of Pronoun Shift results. Each violin plot shows the distribution of model-level sensitivity scores across all evaluated attributes. The black line indicates the mean sensitivity score, with vertical bars denoting the 95% confidence interval. Wider sections of the violin reflect more frequent sensitivity values.
@@ 4.3 Prompt Sensitivity Evaluation
APD ranges from 0 (identical distributions) to 1 (fully divergent), meaning that the score is zero when the model output distribution doesn’t change between the prompt conditions and one if this change is maximal. This serves as the basis for the two sensitivity scores: the Gender Salience Effect Score and the Instruction Presence Effect Score. The two only vary in the prompt condition we’re summing over.
@@
We compute the sensitivity of the tested LLMs to the prompt conditions in three stages: (1) We compute the Absolute Proportion Difference between matched prompt variants (e.g., $\text{Gender}^{+}\text{Instr}^{+}$ vs. $\text{Gender}^{+}\text{Instr}^{-}$ ) at the attribute level. (2) We average Absolute Proportion Difference Scores across all attributes to obtain Gender Salience Effect and Instruction Presence Effect per task. (3) We average GenEffect and InstrEffect Scores across tasks, representing overall sensitivity to prompt structure and refer to this as the Probability of Pronoun Shift.
@@ 5.1 Pronoun Choices Shift when Gender Bias Evaluation is Salient
Figure 1 shows the results for computing the Probability of Pronoun Shift (as defined in Section 4.3) for all models, i.e., the distribution of the prompt sensitivity scores across models. If models were insensitive to the prompt conditions, they would assign the same pronoun in a given scenario, resulting in a sensitivity score of 0. If they show maximally distinct behavior in their gender assignment across conditions, the sensitivity score would be 1.
@@
We observe distinct patterns in how individual models respond to each framing dimension. In the Gender Salience condition (Figure 2a), most models exhibit moderate to high sensitivity, with average scores largely around 0.6. This indicates that when the gender-inference nature of the task is made explicit, models frequently adjust their pronoun outputs. A notable exception is Meta-Llama-3.1-8B, which shows low sensitivity when the prompt primes for gender-related concepts, suggesting a relative insensitivity compared to the other models. This effect appears to drive that Meta-Llama-3.1-8B is an outlier in the overall pronoun shift results (Figure 1).
@@
Figure 3: Pronoun-Specific Shift Probabilities across Models. Each bar represents the mean shift in token probability for a given pronoun (he, she, or they) across all prompt conditions and attributes. The shift is computed using relative differences in pronoun probabilities between paired prompt conditions, rather than absolute differences. Error bars indicate 95% confidence intervals, showing variability across attributes.
@@ 5.3 Effects on the Pronoun-Level
While the previous results indicate that using prompts with common bias evaluation setups change model behavior, it leaves open whether these changes in model behavior are interpretable. Prior work has shown that LLMs often default to assigning male gender when the context is ambiguous Kotek et al. (2023); Dong et al. (2024); Kaneko et al. (2024); Tang et al. (2024). Based on the reasoning that LLMs might learn fairer behavior particularly in evaluation settings, we predict that generally underrepresented genders (“she”) and neutral pronouns (singular “they”) should show an increase in assignment in testing scenarios, in contrast to generally overrepresented genders (“he”). The results are shown in  Figure 3. Pronouns that have a sensitivity of zero don’t change in distribution with varying prompts. Pronouns with a sensitivity < 0 are likely to disappear when the prompt saliently signals gender evaluation and pronouns with a sensitivity > 0 are assigned higher preference.
@@ 5.4 Effects on the Metric-Level
To ensure interpretable data for the proportion of choice analysis, we filter out task-condition pairs in which the model consistently fails to generate pronouns. Specifically, if more than 60% of outputs in a given (model, task, condition) combination contain non-pronoun completions, the setting is considered over-capacity and excluded from analysis. Phi-3-small-128k-instruct and Mistral-Small-Instruct-2409, the smallest models in our evaluation, exhibit the highest number of exclusions, suggesting that potentially limited capacity may impair their ability to provide relevant responses (i.e., outputs containing gendered pronouns or descriptors). Expectedly, instruction-absent conditions were especially noisy in their output but are sufficiently present across models and tasks to allow for an aggregated analysis. We provide all details on the exclusions in Appendix A.
@@ 6.1 Implications for Benchmark Design
Beyond these observations, our work recommends several directions for more robust evaluation design. Conceptually, this challenge is reminiscent of the problem of task demand characteristics in psychology Banaji and Hardin (1996); Greenwald et al. (1998); Oakhill et al. (2005), where participants adjust their behavior once they infer the purpose of a study. Decades of research has brought forth strategies to minimize task demand in people, and similar approaches may help benchmarks better reveal persistent model Morehouse et al. (2025). For example, future benchmarks could explore the benefit of using filler items, refocus on implicit measures (what are reaction time analyses for human studies), or introduce secondary tasks to reduce the prevalence of evaluation-typical features. At a more practical level, benchmarks should rely on diverse prompt sets rather than isolated items, report sensitivity ranges to highlight framing effects, and avoid instruction-heavy setups unless fairness alignment in such settings is the explicit goal. Together, these practices may help ensure that benchmark outcomes reflect underlying model biases rather than prompt-induced compliance.
6.2 (Absence of) Implications for Evaluation Awareness Theory
While the previous paragraph draws parallels to people’s task demand characteristics, we also emphasize an important distinction: While task demand characteristics appear to be driven by people’s meta-awareness about the tasks, this is not a claim we feel positioned to make for models based on the above evidence.
Our results are in line with a rich body of work that has highlighted how characteristics about the data may lead models to learn undesired features (see, e.g., Duchi and Namkoong, 2021; Creager et al., 2021; Gururangan et al., 2018). For example, Language and Vision Models have been shown to leverage spurious correlations in the training data to achieve increased test time performance (most famously, Vision Models categorizing birds based on water vs. land in the background instead of bird-specific features Sagawa et al. (2019); Izmailov et al. (2022)). Our findings can be reframed in terms of any of these phenomena being purely attributable to statistical learning and optimization characteristics.
In this way, our work is very distinct from recent lines of work that investigate models’ meta-awareness Laine et al. (2024); Meinke et al. (2024); Greenblatt et al. (2024), specifically their evaluation awareness and attributing “intentionality” for changing response patterns Needham et al. (2025). While our results are compatible with such a theory, we think they also follow from the simpler assumption that there is a distributional difference between common bias-testing scenarios seen during training and the non-bias-related training distribution. We do not see a need to additionally posit “evaluation awareness” to explain these results.
6.3 Intervention: Using Evaluation Distribution Shifts for Desired Outcomes
Our results have further implications for prompt design as intervention. Prompts that foreground gender concepts can shift model outputs in ways that align with fairness goals. This suggests that strategically framed prompts could serve as lightweight mechanisms to influence LLM behavior in practice—though we must be careful not to mistake prompt compliance for true debiasing.
@@ 6.3 Intervention: Using Evaluation Distribution Shifts for Desired Outcomes
We validate the promise of practically incorporating these insights using two recently proposed gender bias benchmarks Dong et al. (2024); Onorati et al. (2023). After replicating their findings, we adapted their prompts to increase the salience of the gender testing variable. (We report all data and implementational details in Appendix B.) In line with our previous results, we find that for both benchmarks and across tested models gender bias scores significantly shift, sometimes even reversing the previously attested bias trend (see Figure 5 for a summary of the main results). These results emphasize the brittle nature of prompt-based model behavior overall and how gender associations within the task can fundamentally alter gender bias behavior—maybe sometimes even for the better.
@@
Figure 5: Pronoun and Bias Score Shift in bias benchmark replication and intervention conditions. (a) shows overall Probability of Pronoun Shift; (b) shows change in Gender Salience Effect Score). Both reflect aggregate results across all models after prompt modifications. The red line and arrow in (b) indicate the direction and magnitude of the Gender Salience Effect change.